I need to name the slope and name one point on the line that each point-slope equation represents. This is the problem: y = 3+4(x-5)
Answer by kkasko(46)

I need to name the slope and name one point on the line that each point-slope equation represents. This is the problem: y = 3+4(x-5)
Y=3+4(x-5)
Y=3+4x-20
Y=4/1x-17
Slope= 4/1x or 4x and the point on the line or b that each equation represents is (0,-17)